# Large Language Models for Autonomous Driving (LLM4AD): Concept, Benchmark, Experiments, and Challenges

### v3

@@ Abstract
With the broader usage and highly successful development of Large Language Models (LLMs), there has been a growth of interest and demand for applying LLMs to autonomous driving technology. Driven by their natural language understanding and reasoning ability, LLMs have the potential to enhance various aspects of autonomous driving systems, from perception and scene understanding to language interaction and decision-making. In this paper, we first introduce the novel concept of designing LLMs for autonomous driving (LLM4AD). Then, we propose a comprehensive benchmark for evaluating the instruction-following abilities of LLM4AD in simulation. Furthermore, we conduct a series of experiments on real-world vehicle platforms, thoroughly evaluating the performance and potential of our LLM4AD systems. Finally, we envision the main challenges of LLM4AD, including latency, deployment, security and privacy, safety, trust and transparency, and personalization. Our research highlights the significant potential of LLMs to enhance various aspects of autonomous vehicle technology, from perception and scene understanding to language interaction and decision-making.
@@ I Introduction
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=10em,font=,, where level=2text width=11em,font=,, where level=3text width=18em,font=,, [ Large Language Models for Autonomous Driving, ver [ Concept of LLM4AD ] [ Benchmark and Simulation [ LaMPilot-Bench, leaf ] [ CARLA Leaderboard 1.0 Benchmark, leaf ] [ Baselines and Results, leaf ] [ Summary, leaf ] ] [ Real-Vehicle Experiments:
On-Cloud LLM for Personalized
Decision-Making [ Talk2Drive Framework, leaf ] [ Autonomous Driving System, leaf ] [ Experiment Setup, leaf ] [ Experiment Results, leaf ] ] [ Real-Vehicle Experiments:
On-Edge VLM for Personalized
Motion Control [ Problem Statement, leaf ] [ Our Framework, leaf ] [ Real-World Experiment, leaf ] ] [ Challenges [ Latency, leaf ] [ Security and Privacy, leaf ] [ Development, leaf ] [ Safety, leaf ] [ Personalization, leaf ] [ Trust and Transparency, leaf ] ] ]
@@ I-A Background
The recent development of Large Language Models (LLMs) has enhanced numerous cutting-edge technologies, introducing significant advancements across a broad spectrum of applications [zhao2023survey, wang2023survey]. Their applications span from the classic nature language process (NLP) tasks like document modification and information extraction to the new emerging scenarios such as LLM-based agents and LLMs for evaluation [zhao2023survey]. One particular application that our study focuses on, among many other domains, is adopting LLMs for autonomous driving (LLM4AD). In this area, various advanced emerging LLM-based algorithms and technologies are continually enhancing the capabilities of autonomous driving technology, leveraging the potential of LLMs to drive innovation and efficiency forward. LLMs can contribute to autonomous systems from high-level decision-making processes to meticulous low-level control. On the high-level side, LLMs can actively engage in adjusting driving mode or decision-making process [cui_drive_2024]. Imagine the scenario in which you are seated in an autonomous vehicle, where all you need to do is abstractly express your feelings like “I do not want my friends to wait for me.” The LLM4AD system then interprets your emotions and adjusts the vehicle’s control strategies accordingly, aligning with your current driving mood or preferences. By contrast, non-LLM-based systems lack the ability to accurately comprehend or interpret humans’ intentions only from some vague expressions [cui2023large]. Besides, LLMs-based systems have significant potential for achieving personalization in autonomous driving due to their continuous learning capacity. This capability enables it to continuously adapt to individuals’ preferences and desires, improving the driving experience according to different users. In addition, LLMs also hold the potential to develop a knowledge-driven system capable of driving like an expert, which can accumulate experience through continuous driving [fu_drive_2024].
@@
  • Intuitive Language Interaction: LLMs enable intuitive communication between humans and vehicles. Humans can express abstract commands and feelings, while the LLMs can accurately capture the intentions behind human expressions.
@@ I-B Overview of This Article
Fig. 1 provides an overview of the reminder structure of this paper. Our paper presents a comprehensive study on integrating LLMs into autonomous driving systems, from the proposed concept and the creation of an open dataset to the implementation in simulations and real-vehicle experiments. Sec. II introduces the concept and key elements of our proposed LLM4AD framework. Sec. III develops an open benchmark designed to evaluate the performance of LLMs in various autonomous driving tasks, which includes a dataset, evaluation metrics, and baseline models. Sec. IV and Sec. V present the implementation of our framework on a real-vehicle platform and discuss the results of a series of experiments conducted to validate the system’s driving and personalization performance. Sec. VI analyzes the current challenges and outlines promising future research directions for leveraging LLMs in autonomous driving applications. Finally, Sec. VII summarizes the key contributions and findings of this work.
@@
TABLE I: Symbols and their meanings used in this document.
Symbol Meaning
LLM Large Language Model
$I$ Instruction
$C$ Contextual (Environment) Information
$S$ System Message
$H$ History Data
$F$ Feedback
$R$ Reasoning Thoughts
$P$ Generated Program/Policy
$V$ Visual Input
$D$ Scene Description
$A$ Executed Action
LMP Language Model Program
$\ell$ LLM backbone
API Application Programming Interface
TTC Time to Collision
SV Speed Variation
TE Time Efficiency
@@ II Concept of LLM4AD
In this section, we introduce a perspective where LLMs play the role of the decision-making “brain” within the autonomous driving system. Within our framework, LLMs do not directly affect perception or localization modules, which serve as the vehicle’s “eyes;” instead, they utilize outputs from these modules as references to guide the high-level decision-making process. By receiving processed data from them, LLMs can enhance informed decision-making, leading to significant improvements in the performance of autonomous vehicles. At the downstream, the vehicle’s control module serves as its “hands,” executing the driving policy obtained from the LLM-based decision-making process. The symbols we used in the rest of the paper are listed in Tab. I.
The overall LLM4AD framework is shown in Fig. 2. In our framework, humans provide instructions and evaluations, where the instructions $I$ and evaluations $F$ along with historical memories $H$, system messages $S$, and context information $C$ serve as inputs to LLMs. The memory module stores corresponding historical interactions between humans and vehicles $H$ for different users. Upon receiving these inputs, LLMs engage in reasoning and produce outputs, including generated Language Model Programs (LMPs) $P$ and reasoning thoughts $R$. The generated LMP is sent to the executor for execution in the environment, while the reasoning thoughts help LLMs generate more reasonable driving policies. It is important to note that this is a general concept, where specific implementations may vary across different applications. However, all applications follow the principles and framework of this concept.
@@ II-A Human Instruction and Evaluation
Human instructions $I$ and evaluations $F$ are directly inputted to LLMs by humans in natural language. $I$ includes the humans’ desired needs for the autonomous agent, while human evaluations $F$ are the feedback on the effectiveness of these driving policies.
@@ II-B System Message
The system message $S$ is a way to provide instructions or context to the LLM4AD system at the start of a conversation or task. System messages $S$ allow developers to clearly define the role, constraints, and objectives that LLM4AD systems should follow when engaging with users or working on particular types of tasks.
@@ II-D History Memory and Memory Module
The memory module stores profiles from various users to enhance driving experiences for all users. Whenever a human user utilizes the LLM4AD system, the system logs the relevant historical interactions $H$ associated with that user. Subsequently, the historical data for the current human user is transmitted to LLMs as input. This particular history interaction $H$ serves as a reference point for the preferences of the current users, therefore guiding the system to improve user experience. After each trip, the interaction data will be updated in the corresponding profile in the memory module. The specifics of the interaction data vary for each system and will be elaborated in their respective sections.
@@ II-E Large Language Models
LLMs serve as the core module in our framework, which will receive all the inputs (history interactions) $H$, system message $S$, situation description $C$, human instruction $I$, and human evaluation $F$) and generate the textual outputs (LMPs $P$ and reasoning thoughts $R$). One important note is that we employ the chain-of-thought prompting [wei_chain--thought_2022] technique, which serves as a guiding signal to ensure alignment with human-like reasoning and practical driving considerations.
@@ II-F Generated Policy/Program
Inspired by the concept of “Code as Policy [liang_code_2023],” one of the key outputs from LLMs is the generated LMPs $P$ consisting of executable codes. These codes are used to affect the driving behavior of the ego agent within the environment, and they not only have the ability to generalize to new natural language commands but can also provide precise numerical values such as velocities based on vague linguistic descriptions (e.g., “hurry,” “turn left”) depending on the driving context.
II-G Output Thought
By employing chain-of-thought prompting [wei_chain--thought_2022], LLMs not only generate the program code but also provide a step-by-step explanation of the thought process used to reach that solution. These chains of thoughts represent LLMs’ reasoning behind each decision, such as “Since the command is “hurry,” I will increase the target velocity” or “To turn left, I need to adjust the steering angle based on the current speed.” The output thoughts $R$ accompany the generated programs $P$, offering insights into how LLMs interpret natural language commands in the context of the driving situation description to produce precise control values like velocities or steering angles. This output thought improves the transparency and interpretability of the LLM4AD system’s decision-making process.

### v4

# LLM4AD: Large Language Models for Autonomous Driving — Concept, Review, Benchmark, Experiments, and Future Trends

@@ Abstract
With the broader adoption and highly successful development of Large Language Models (LLMs), there has been growing interest and demand for applying LLMs to autonomous driving technology. Driven by their natural language understanding and reasoning capabilities, LLMs have the potential to enhance various aspects of autonomous driving systems, from perception and scene understanding to interactive decision-making. In this paper, we first introduce the novel concept of designing Large Language Models for Autonomous Driving (LLM4AD), followed by a review of existing LLM4AD studies. Then, we propose a comprehensive benchmark for evaluating the instruction-following and reasoning abilities of LLM4AD systems, which includes LaMPilot-Bench, CARLA Leaderboard 1.0 Benchmark in simulation and NuPlanQA for multi-view visual question answering. Furthermore, we conduct extensive real-world experiments on autonomous vehicle platforms, examining both on-cloud and on-edge LLM deployment for personalized decision-making and motion control. Next, we explore the future trends of integrating language diffusion models into autonomous driving, exemplified by the proposed ViLaD (Vision-Language Diffusion) framework. Finally, we discuss the main challenges of LLM4AD, including latency, deployment, security and privacy, safety, trust and transparency, and personalization.
@@ I Introduction
Large Language Models for Autonomous Driving
Concept of LLM4AD Review of Existing LLM4AD Studies Learning Method Pipeline Evaluation Deployment Platform Benchmark and Simulation LaMPilot-Bench CARLA Leaderboard 1.0 Benchmark NuplanQA Benchmark Summary Real-Vehicle Experiments: On-Cloud LLM for Personalized Decision-Making Talk2Drive Framework Autonomous Driving System Experiment Setup Experiment Results Real-Vehicle Experiments: On-Edge VLM for Personalized Motion Control Problem Statement Our Framework Real-World Experiment Future Trend of LLM4AD: Language Diffusion Models in Autonomous Driving ViLaD: Vision Language Diffusion Experiment and Results Real-World Case Studies Challenges Latency Security and Privacy Development Safety Personalization Trust and Transparency
@@ I-A Background
The recent development of Large Language Models (LLMs) has enhanced numerous cutting-edge technologies, introducing significant advancements across a broad spectrum of applications [zhao2023survey, wang2023survey]. Their applications span from the classic natural language processing (NLP) tasks like document modification and information extraction to the new emerging scenarios such as LLM-based agents and LLMs for evaluation [zhao2023survey]. One particular application that our study focuses on, like many other domains, is adopting LLMs for autonomous driving (LLM4AD). In this area, various advanced emerging LLM-based algorithms and technologies are continually enhancing the capabilities of autonomous driving technology, leveraging the potential of LLMs to drive innovation and efficiency forward. LLMs can contribute to autonomous systems from high-level decision-making processes to meticulous low-level control.
On the high-level side, LLMs can actively engage in adjusting driving mode or decision-making process [cui_drive_2024]. Imagine the scenario in which you are seated in an autonomous vehicle, where all you need to do is to abstractly express your feelings like “I do not want my friends to wait for me.” The LLM4AD system then interprets your emotions and adjusts the vehicle’s control strategies accordingly, aligning with your current driving mood or preferences. By contrast, non-LLM-based systems lack the ability to accurately comprehend or interpret humans’ intentions only from some vague expressions [cui2023large]. Besides, LLM-based systems have significant potential for achieving personalization in autonomous driving due to their continuous learning capability. This capability enables it to continuously adapt to individuals’ preferences and desires, improving the driving experience according to different users. In addition, LLMs also hold the potential to develop a knowledge-driven system capable of driving like an expert, which can accumulate experience through continuous driving [fu_drive_2024].
@@
Besides applying LLMs into the conventional module-based autonomous systems, LLMs also play a significant role in the end-to-end system for autonomous driving. By leveraging their powerful reasoning and multimodal understanding capabilities, LLMs can directly process sensory data and ego-vehicle states to generate planning trajectories and fine-grained control actions [chen2024end]. This LLM-driven paradigm has attracted increasing attention from both industry [driveirl, hu2023imitation, pmlr-v164-scheel22a] and academia [hu_planning-oriented_2023]. Unlike the conventional modular-based paradigms that separately handle perception, prediction, planning, and control, the LLM-based end-to-end autonomous driving system is fully differentiable and optimizes the entire driving process simultaneously.
@@
  • Intuitive Language Interaction: LLMs enable intuitive communication between humans and vehicles. Humans can express abstract commands and feelings to the vehicle, while the LLMs can accurately capture the intentions behind human expressions.
@@ I-B Overview of This Article
Fig. 1 provides an overview of the structure of this paper. Sec. II introduces the concept and key elements of our proposed LLM4AD framework. Sec. III reviews existing works related to the integration of LLMs into autonomous driving systems. Sec. IV develops open benchmarks designed to evaluate the performance of LLMs in various autonomous driving tasks, which includes datasets, evaluation metrics, and baseline models. Sec. V and Sec. VI present the implementation of our framework on a real-vehicle platform and discuss the results of a series of experiments conducted to validate the system’s driving and personalization performance. Sec. VII proposes the future trend of leveraging large language diffusion models in autonomous driving. Sec. VIII analyzes the current challenges and outlines promising research directions for leveraging LLMs in autonomous driving applications. Finally, Sec. IX summarizes the key contributions and findings of this work.
@@ II Concept of LLM4AD
In this section, we introduce a perspective where LLMs play the role of the decision-making “brain” within the autonomous driving system. Within our framework, LLMs do not directly affect perception or localization modules, which serve as the vehicle’s “eyes;” instead, they utilize outputs from these modules as references to guide the high-level decision-making process. By receiving processed data from them, LLMs can enhance informed decision-making, leading to significant improvements in the performance of autonomous vehicles. At the downstream, the vehicle’s control module serves as its “hands,” executing the driving policy obtained from the LLM-based decision-making process.
The overall LLM4AD framework is shown in Fig. 2. In our framework, humans provide instructions and evaluations, where the instructions $I$ and feedback/evaluations $F$ along with historical memories $H$, system messages $S$, and context information $C$ serve as inputs to LLMs. The memory module stores corresponding historical interactions between humans and vehicles $H$ for different users. Upon receiving these inputs, LLMs engage in reasoning and produce outputs, including generated Language Model Programs (LMPs) or policies $P$ and reasoning thoughts $R$. The $P$ is then sent to the executor for execution in the environment, while the reasoning thoughts help LLMs generate more reasonable driving policies. It is important to note that this is a general concept, where specific implementations may vary across different applications. However, all applications follow the principles and framework of this concept.
@@ II-A Human Instruction and Evaluation
Human instructions $I$ and evaluations $F$ are directly fed into LLMs in natural language. $I$ includes the humans’ desired needs for the autonomous agent, while human evaluation $F$ is the feedback on the effectiveness of these driving policies.
@@ II-B System Message
The system messages $S$ provide instructions or context to the LLM4AD system at the start of a conversation or task. System messages $S$ allow developers to clearly define the role, constraints, and objectives that LLM4AD systems should follow when engaging with users or working on particular types of tasks.
@@ II-D History Memory and Memory Module
The memory module stores profiles from various users to enhance driving experiences for all users. Whenever a human user utilizes the LLM4AD system, the system logs the relevant historical interactions $H$ associated with that user. Subsequently, the historical data for the current human user is transmitted to LLMs as input. This particular historical interaction $H$ serves as a reference point for the preferences of the current users, therefore guiding the system to improve user experience. After each trip, the interaction data will be updated in the corresponding profile in the memory module. The specifics of the interaction data vary for each system and will be elaborated in their respective sections.
@@ II-E Large Language Models
LLMs serve as the core module in our framework, which will receive all the inputs (history interactions) $H$, system message $S$, situation description $C$, human instruction $I$, and human feedback $F$ and generate the textual outputs (Generated Policies/LMPs $P$ and reasoning thoughts $R$). One important note is that we employ the chain-of-thought prompting [wei_chain--thought_2022] technique, which serves as a guiding signal to ensure alignment with human-like reasoning and practical driving considerations.
@@ II-F Generated Policy/Program
Inspired by the concept of “Code as Policy [liang_code_2023],” one of the key outputs from LLMs is the generated Policies/s $P$ consisting of executable codes. These codes are used to affect the driving behavior of the ego agent within the environment, and they not only have the ability to generalize to new natural language commands but can also provide precise numerical values such as velocities based on vague linguistic descriptions (e.g., “hurry,” “turn left”) depending on the driving context.
II-G Output Thoughts
By employing chain-of-thought prompting [wei_chain--thought_2022], LLMs not only generate the program code but also provide a step-by-step explanation of the thought process used to reach that solution. These chains of thoughts represent LLMs’ reasoning behind each decision, such as “Since the command is ‘hurry,’ I will increase the target velocity” or “To turn left, I need to adjust the steering angle based on the current speed.” The output thoughts $R$ accompany the generated policies/LMPs $P$, offering insights into how LLMs interpret natural language commands in the context of the driving situation description to produce precise control values like velocities or steering angles. This output thought improves the transparency and interpretability of the LLM4AD system’s decision-making process.